TY 2 1. Calculate the molarity of (mol/L) of 3% hydrogen peroxide and record it in Data Table 2. The term 3% concentration
I dont need #5. I’ve got that one. I’m just confused on the others.
TY 2 1. Calculate the molarity of (mol/L) of 3% hydrogen peroxide and record it in Data Table 2. The term "3% concentration" is a weight-to-weight measurement, which means that there are 3 g solute per 100 g solvent. The molar mass of hydrogen peroxide is 34.01 g/mol. 2. Calculate the moles of hydrogen peroxide present in each trial. 3. Calculate the moles of oxygen generated in each trial. 4. Convert the change in volume from Activity1 to liters. 5. Convert the air temperature from Activity 1 from Celsius to Kelvin.
Data Table 1 Trial 8 Trial 6 Trial7 (3 mL (4 mL H2O2) H2O2) Trial 1 Trial 2 Trial 3 Trial 4 Trial 5 (1 mL (1 mL (2 mL (2 mL (3 mL H2O2) H2O2) H2O2) H2O2) H2O2) (4 mL H2O2) Air 24 24 24 24 23 temperature °C) Volume H2O2 liquid (mL) 4 4 3 2 2 1 Initial Volume Gas (mL) 13mL 14mL 10mL 19mL 10mL 18mL 26mL 21mL Final Volume 39mL >50mL >50mL 19mL 47mL 23mL 20mL 36mL Gas (mL) 28mL 36mL 37mL AV (mL) 10mL 9mL 29mL 2mL 2mL ch 24 24 23
Answer #1

Concentration : 3 % w/w

Molar mass of H2O2 = 34.01 g/mol

Let us consider density of solvent(water) = 1g/ml

So, 100 g of solvent = 100 ml of water = 0.1 lt of solvent

Molarity = (No. of moles/Vol in lts) = (Mass/molar mass)/Vol in lts = (3 g/34.01 g/mol)/0.1 lt = 0.882 M

No. of moles of H2O2 in each trials :

Trial 1 and Trial 2 : 1 ml = 0.001 lts = Molarity x vol in lts = 0.882 M x 0.001 lts = 0.000882 moles

Trial 3 and trial 4 : 2 ml = 0.002 lts = 0.882 M x 0.002 lts = 0.001764 moles

Trial 5 and Trial 6 : 3 ml = 0.003 lts = 0.882 M x 0.003 lts = 0.002646 moles

Trial 7 and trial 8 : 4 ml = 0.004 lts = 0.882 M x 0.004 lts = 0.003528 moles

   H2O2 (aq) ----> H2O(l)+ 1/2O2 (g)

Molar ratio of H2O2 and O2 = 2:1

Theoretical moles of oxygen in each trails :

Trial 1 and Trial 2 : 1 ml = 0.001 lts = 0.000882 /2 = 0.000441 moles

Trial 3 and trial 4 : 2 ml = 0.002 lts = 0.001764/2 = 0.000882 moles

Trial 5 and Trial 6 : 3 ml = 0.003 lts = 0.002646/2 = 0.001323 moles

Trial 7 and trial 8 : 4 ml = 0.004 lts = 0.003528/2 = 0.001764 moles

Change in volume in lts for each trial :

Trial 1 and trial 2 : 2 ml = 2/1000 = 0.002 lts

Trial 3 : 10 ml = 10/1000 = 0.01 lts

Trial 4 : 9 ml = 9 /1000 = 0.009 lts

Trial 5 : 28 ml = 28/1000 = 0.028 lts

Trial 6 : 29 ml = 29/1000 = 0.029 lts

Trial 7 : 36 ml = 36/1000 = 0.036 lts

Trial 8 : 37 ml = 37/1000 = 0.037 lts